The weights of newborn baby boys born at a local hospital are believed to have a normal distribution with a mean weight of 3331

3331 grams and a variance of 362,404

362,404. If a newborn baby boy born at the local hospital is randomly selected, find the probability that the weight will be less than 3812

3812 grams. Round your answer to four decimal places.


1
Expert's answer
2022-03-21T00:30:18-0400

"P(X<3812)=P(Z<\frac{3812-3331}{362.404})=P(Z<1.33)=0.9082."